# Long distance propagation of light in random media with partially coherent sources

### v1

@@ 6.1 Limiting moments in the kinetic regime
$\lim_{\varepsilon\to 0}\mathcal{R}(\eta(y_{\gamma_{2}}-x_{\gamma_{1}}),\eta\varepsilon^{\beta-1}(\xi_{\gamma_{1}}-\zeta_{\gamma_{2}}))=\begin{cases}\mathcal{R}(y_{\gamma_{2}}-x_{\gamma_{1}},0),\quad\beta>1\\ \mathcal{R}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}}),\quad\beta=1\,.\end{cases}$
@@
$M_{p,p}(z,r,X,Y;T)$ $=S_{p}(r,\ldots,r;T)e^{-\frac{pk_{0}^{2}R(0)z}{4}}\big{(}1+\sum\limits_{\kappa=1}^{K}\prod\limits_{\gamma\in\Lambda_{\kappa}}[\mathcal{R}(y_{\gamma_{2}}-x_{\gamma_{1}},0)-1]\big{)}$ (65)
@@
$M_{1,1}^{\infty}(z,r,x,y)=\mathcal{Q}(y-x,0)\,.$ (66)
for $\mathcal{Q}$ as in (32).
@@
$+e^{-\frac{pk_{0}^{2}R(0)z}{4}}\sum\limits_{\kappa=1}^{K}\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}(\xi_{j}-\zeta_{j})\cdot r}\prod\limits_{\gamma\in\Lambda_{\kappa}}[\mathcal{R}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})-1]\hat{S_{p}}(v;T)\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}\,.$
@@
×$\times\int\limits_{\mathbb{R}^{pd}}\prod\limits_{\gamma\in\Lambda_{\kappa}}\delta_{0}(x^{\prime}_{\gamma_{1}}-y^{\prime}_{\gamma_{2}})[\mathcal{R}(y_{\gamma_{2}}-x_{\gamma_{1}},\zeta_{\gamma_{2}})-1]e^{i\zeta_{\gamma_{2}}\cdot(r-y^{\prime}_{\gamma_{2}})}\prod\limits_{j\notin\Lambda_{\kappa,1}}\delta_{0}(r-x^{\prime}_{j})\prod\limits_{l\notin\Lambda_{\kappa,2}}\delta_{0}(r-y^{\prime}_{l})\prod_{j=1}^{p}\frac{\mathrm{d}\zeta_{j}}{(2\pi)^{d}}\,.$
@@
$\overline{M}^{\infty}_{1,1}(z,r,x,y)=e^{-\frac{k_{0}^{2}R(0)z}{4}}\int\limits_{\mathbb{R}^{d}}[\mathcal{R}(y-x,\zeta)-1]e^{i\zeta\cdot r}\frac{\mathrm{d}\zeta}{(2\pi)^{d}}\,.$ (69)
@@
In the case $\beta=1$ and $\theta\ll 1$, we still have
$M^{\theta}_{p,p}(z,r,X,Y;T)$ $:=\lim_{\varepsilon\to 0}m^{\varepsilon}_{p,p}(z,r,X,Y;T)=e^{-\frac{pk_{0}^{2}R(0)z}{4}}\Big{(}S_{p}(r,\ldots,r;T)$
$+\sum\limits_{\kappa=1}^{K}\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}(\xi_{j}-\zeta_{j})\cdot r}\prod\limits_{\gamma\in\Lambda_{\kappa}}[\mathcal{R}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})-1]\hat{S_{p}}(v;T)\prod\limits_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}\,.$
$I^{\theta}_{0}:=\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}(\xi_{j}-\zeta_{j})\cdot r}\prod\limits_{\gamma\in\Lambda_{\kappa}}[\mathcal{R}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})-1]\hat{S_{p}}(v;T)\mathrm{d}\tilde{\Xi}$
$=\theta^{pd}\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}(\xi_{j}-\zeta_{j})\cdot r}\sum\limits_{\pi_{p}}F(t_{j}-t^{\prime}_{\pi_{p}(j)})\hat{\Gamma}\Big{(}\xi_{j}-\zeta_{\pi_{p}(j)},\frac{\theta(\xi_{j}+\zeta_{\pi_{p}(j)})}{2}\Big{)}\prod\limits_{\gamma\in\Lambda_{\kappa}}[\mathcal{R}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})-1]\mathrm{d}\tilde{\Xi}.$
$I^{\theta}_{0}=\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}\zeta_{j}\cdot r}\sum\limits_{\pi_{p}}\prod\limits_{j=1}^{p}F(t_{j}-t^{\prime}_{\pi_{p}(j)})\hat{\Gamma}(\zeta_{\pi_{p}}(j),\xi_{j})\prod\limits_{\gamma\in\Lambda_{\kappa}}[\mathcal{R}\Big{(}y_{\gamma_{2}}-x_{\gamma_{1}},\frac{\xi_{\gamma_{1}}-\xi_{\pi_{p}^{-1}(\gamma_{2})}}{\theta}+\frac{\zeta_{\pi_{p}(\gamma_{1})}+\zeta_{\gamma_{2}}}{2}\Big{)}-1]\mathrm{d}\tilde{\Xi}.$
$\lim_{\theta\to 0}\mathcal{R}\Big{(}y_{\gamma_{2}}-x_{\gamma_{1}},\frac{\xi_{\gamma_{1}}-\xi_{\pi_{p}^{-1}(\gamma_{2})}}{\theta}+\frac{\zeta_{\pi_{p}(\gamma_{1})}+\zeta_{\gamma_{2}}}{2}\Big{)}=1\,.$
$&M_{p,p}(z,r,X,Y;T)=e^{-\frac{pk_{0}^{2}R(0)z}{4}}\Big{(}S_{p}(r,\ldots,r;T)\\ &+\sum\limits_{\kappa=1}^{K}\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}M_{1,1}(z,r,x_{\gamma_{1}},y_{\gamma_{2}};t_{\gamma_{1}},t^{\prime}_{\gamma_{2}})-M_{0,0}(z,r;t_{\gamma_{1}},t^{\prime}_{\gamma_{2}})\Big{)}\sum\limits_{\pi_{\bar{m}(\kappa)}}\prod\limits_{j=1}^{\bar{m}(\kappa)}M_{0,0}(z,r;t_{s_{j}},t^{\prime}_{s^{\prime}_{\pi_{\bar{m}(\kappa)}}(j)})\,,\end{aligned}$ (71)
@@ 6.2 Limiting moments in the diffusive regime
$\lim_{\varepsilon\to 0}U_{\eta}^{p}\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}\mathcal{R}\big{(}\eta(y_{\gamma_{2}}-x_{\gamma_{1}}),\eta\varepsilon^{\beta-1}(\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})\big{)}-1\Big{)}$
@@ 8 Conclusions and possible extensions
The goal of this paper was to provide a complete statistical description of wavefields due to partially coherent incident beams under the scintillation scaling of the Itô-Schrödinger white noise paraxial regime. We observe that statistical fluctuations in the time averaged intensity are amplified or reduced depending on the coherence width of the source relative to the fluctuations in the medium and the averaging time at the detector. In particular, this leads to scintillation indices ranging from 3 to 0 in the diffusive limit. The results presented in this paper can potentially be useful in providing a theoretical guidance to optimize performance of communication links. The diffusive regime considered here provides a simplified mapping between the turbulence parameters of the medium, source statistics and intensity measurements. The asymptotic statistical behaviour of wavefields in this regime is primarily governed by the macroscopic scaling parameters rather than the intricate profiles of the incident beams themselves. While the kinetic regime does not immediately lead to as simplified models as in the diffusive regime, it may prove useful for engineering applications when complemented with numerical simulations. Quantifying the range of validity of these regimes and further numerical experiments in the kinetic regime are left for future investigation.
@@ Appendix A Operator $L^{\varepsilon}_{p,q}$
For $1\leq j<j^{\prime}\leq p$ and $p+1\leq j<j^{\prime}\leq p+q$, we define $B_{j,j^{\prime}}$ as a $d(p+q)\times d$ matrix of $d\times d$ blocks with block $j$ as $\mathbb{I}_{d}$ and block $j^{\prime}$ as $-\mathbb{I}_{d}$.

### v2

@@ 6.1 Limiting moments in the kinetic regime
[…]
@@
$M_{p,p}(z,r,X,Y;T)$ $=S_{p}(r,\ldots,r;T)e^{-\frac{pk_{0}^{2}R(0)z}{4}}\big{(}1+\sum\limits_{\kappa=1}^{K}\prod\limits_{\gamma\in\Lambda_{\kappa}}[{\color[rgb]{0,0,0}\mathcal{R}_{z}}(y_{\gamma_{2}}-x_{\gamma_{1}},0)-1]\big{)}$ (65)
@@
$M_{1,1}^{\infty}(z,r,x,y)={\color[rgb]{0,0,0}\mathcal{Q}_{z}}(y-x,0)\,.$ (66)
for $\color[rgb]{0,0,0}\mathcal{Q}_{z}$} as in (32).
@@
$+e^{-\frac{pk_{0}^{2}R(0)z}{4}}\sum\limits_{\kappa=1}^{K}\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}(\xi_{j}-\zeta_{j})\cdot r}\prod\limits_{\gamma\in\Lambda_{\kappa}}[{\color[rgb]{0,0,0}\mathcal{R}_{z}}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})-1]\hat{S_{p}}(v;T)\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}\,.$
@@
$\times\int\limits_{\mathbb{R}^{pd}}\prod\limits_{\gamma\in\Lambda_{\kappa}}\delta_{0}(x^{\prime}_{\gamma_{1}}-y^{\prime}_{\gamma_{2}})[{\color[rgb]{0,0,0}\mathcal{R}_{z}}(y_{\gamma_{2}}-x_{\gamma_{1}},\zeta_{\gamma_{2}})-1]e^{i\zeta_{\gamma_{2}}\cdot(r-y^{\prime}_{\gamma_{2}})}\prod\limits_{j\notin\Lambda_{\kappa,1}}\delta_{0}(r-x^{\prime}_{j})\prod\limits_{l\notin\Lambda_{\kappa,2}}\delta_{0}(r-y^{\prime}_{l})\prod_{j=1}^{p}\frac{\mathrm{d}\zeta_{j}}{(2\pi)^{d}}\,.$
@@
$\overline{M}^{\infty}_{1,1}(z,r,x,y)=e^{-\frac{k_{0}^{2}R(0)z}{4}}\int\limits_{\mathbb{R}^{d}}[{\color[rgb]{0,0,0}\mathcal{R}_{z}}(y-x,\zeta)-1]e^{i\zeta\cdot r}\frac{\mathrm{d}\zeta}{(2\pi)^{d}}\,.$ (69)
@@
$\begin{aligned} M_{p,p}(z,r,X,Y;T)&=\Big{(}M_{p,p}^{\infty}(z,\cdot,X,Y)\ast S_{p}(\cdot;T)\Big{)}(r,\ldots,r)\,.\end{aligned}$ (70)
In the case $\beta=1$ and $\theta\ll 1$, we still have
$M^{\theta}_{p,p}(z,r,X,Y;T)$ $=\lim_{\varepsilon\to 0}m^{\varepsilon}_{p,p}(z,r,X,Y;T)=e^{-\frac{pk_{0}^{2}R(0)z}{4}}\Big{(}S_{p}(r,\ldots,r;T)$
$+\sum\limits_{\kappa=1}^{K}\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}(\xi_{j}-\zeta_{j})\cdot r}\prod\limits_{\gamma\in\Lambda_{\kappa}}[{\color[rgb]{0,0,0}\mathcal{R}_{z}}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})-1]\hat{S_{p}}(v;T)\prod\limits_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}\,.$
$I^{\theta}_{0}=\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}(\xi_{j}-\zeta_{j})\cdot r}\prod\limits_{\gamma\in\Lambda_{\kappa}}[{\color[rgb]{0,0,0}\mathcal{R}_{z}}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})-1]\hat{S_{p}}(v;T)\mathrm{d}\tilde{\Xi}$
$=\theta^{pd}\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}(\xi_{j}-\zeta_{j})\cdot r}\sum\limits_{\pi_{p}}F(t_{j}-t^{\prime}_{\pi_{p}(j)})\hat{\Gamma}\Big{(}\xi_{j}-\zeta_{\pi_{p}(j)},\frac{\theta(\xi_{j}+\zeta_{\pi_{p}(j)})}{2}\Big{)}\prod\limits_{\gamma\in\Lambda_{\kappa}}[{\color[rgb]{0,0,0}\mathcal{R}_{z}}(y_{\gamma_{2}}-x_{\gamma_{1}},\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})-1]\mathrm{d}\tilde{\Xi}$.
$I^{\theta}_{0}=\int\limits_{\mathbb{R}^{2pd}}e^{i\sum_{j=1}^{p}\zeta_{j}\cdot r}\sum\limits_{\pi_{p}}\prod\limits_{j=1}^{p}F(t_{j}-t^{\prime}_{\pi_{p}(j)})\hat{\Gamma}(\zeta_{\pi_{p}}(j),\xi_{j})\prod\limits_{\gamma\in\Lambda_{\kappa}}[{\color[rgb]{0,0,0}\mathcal{R}_{z}}\Big{(}y_{\gamma_{2}}-x_{\gamma_{1}},\frac{\xi_{\gamma_{1}}-\xi_{\pi_{p}^{-1}(\gamma_{2})}}{\theta}+\frac{\zeta_{\pi_{p}(\gamma_{1})}+\zeta_{\gamma_{2}}}{2}\Big{)}-1]\mathrm{d}\tilde{\Xi}$.
$\lim_{\theta\to 0}{\color[rgb]{0,0,0}\mathcal{R}_{z}}\Big{(}y_{\gamma_{2}}-x_{\gamma_{1}},\frac{\xi_{\gamma_{1}}-\xi_{\pi_{p}^{-1}(\gamma_{2})}}{\theta}+\frac{\zeta_{\pi_{p}(\gamma_{1})}+\zeta_{\gamma_{2}}}{2}\Big{)}=1\,.$
@@ 6.2 Limiting moments in the diffusive regime
$\lim_{\varepsilon\to 0}U_{\eta}^{p}\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}{\color[rgb]{0,0,0}\mathcal{R}_{z}}\big{(}\eta(y_{\gamma_{2}}-x_{\gamma_{1}}),\eta\varepsilon^{\beta-1}(\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})\big{)}-1\Big{)}$
@@ 8 Conclusions and possible extensions
The goal of this paper was to provide a complete statistical description of wavefields due to partially coherent incident beams under the scintillation scaling of the Itô-Schrödinger white noise paraxial regime. We observe that statistical fluctuations in the time averaged intensity are amplified or reduced depending on the coherence width of the source relative to the fluctuations in the medium and the averaging time at the detector. In particular, this leads to scintillation indices ranging from 3 to 0 in the diffusive limit. The results presented in this paper can potentially be useful in providing a theoretical guidance to optimize performance of communication links. The diffusive regime considered here provides a simplified mapping between the turbulence parameters of the medium, source statistics and intensity measurements. The asymptotic statistical behaviour of wavefields in this regime is primarily governed by the macroscopic scaling parameters rather than the intricate profiles of the incident beams themselves. While the kinetic regime does not immediately lead to as simplified models as in the diffusive regime, it may prove useful for engineering applications when complemented with numerical simulations. It is also likely that one can arrive at a similar conclusion starting from the paraxial approximation, without explicitly using the white noise model as was done in [6] for coherent beams. Quantifying the range of validity of these regimes and further numerical experiments in the kinetic regime are left for future investigation.
@@ Appendix A Operator $L^{\varepsilon}_{p,q}$
For $1\leq j<j^{\prime}\leq p$ and $p+1\leq j<j^{\prime}\leq p+q$, we define $B_{j,j^{\prime}}$ as a $d(p+q)\times d$ matrix of $d\times d$ blocks with block $j$ as $\mathbb{I}_{d}$ and block $j^{\prime}$ as $\mathbb{I}_{d}$.